OpenStudy (anonymous):

How many total atoms are there in 65.7 g of water (H2O)?

OpenStudy (anonymous):

Okay, so for this, you have to use the avogadro constant 6.02x10^23. You only ever use this number if it mentions atoms or molecules. You have to do a few conversions as well. To use avogadro's constant you have to get your grams to moles. You have 65.7g of water. How many grams of H2O are in one mole of H2O? Well a mole is just the rounded number of the atomic weight. You may need your periodic table. H=1 O=16. There is two hydrogens and one oxygen. So (1x2)+16=18. 18 grams = 1 mole of H2O So far you should have an equation that looks something like this: \[65.7g \times\frac{ 1molH2O }{ 18g H2O }\] Now you can cancel out your grams. 1 mole of whatever you want to figure out will always be the conversion to this crazy avogradro's number. Now you'll have this \[65.7g \times\frac{ 1molH2O }{ 18g H2O } \times \frac{ 6.02\times10^{23} }{ 1molH2O }\] Now you can multiply everything across and divide the sums of each. (65.7x1x6.02)/(18x1) = (395.514)/(18) = [21.973] Then you can slap 10^23 on it. 21.973x10^23 should be your awnser